If an electric wire is allowed to produce a magnetic field no larger than that of the Earth (0.50×10−4T) at a distance of 20 cm from the wire, what is the maximum current the wire can carry?
Answer #1

B = 0.5E-4T

r =20cm =0.2m

to calculate the current, we use the equation: B = \muo I/2\pir

so, I = 2\piBr/\muo = [2\pi * 0.5*10-4 * 0.2]/4\pi*10-7 = 50A